# Semantic Mapping in Indoor Embodied AI – A Survey on Advances, Challenges, and Future Directions

### v2

@@ 1 Introduction
To provide a principled understanding of semantic map-building methods, this survey categorizes the literature along two fundamental axes (see Fig. 1): map structure (e.g., topological graphs, spatial grids, dense geometric, and hybrid representations) and semantic encoding (explicit annotations vs. learned implicit features). This taxonomy reflects core design choices that influence map scalability, interpretability, generalization, multi-modal fusion and querying. By organizing approaches in this manner, we aim to unify diverse strands of research (see Tab. 1 for a summary of the reviewed papers), highlight the trade-offs between different representations and present key challenges and future opportunities in semantic mapping. This survey centers on semantic mapping approaches in the context of indoor mobile robots, a domain that offers a well-defined, practically relevant and technically rich environment for research. We focus on semantic mapping within embodied AI, that enables controlled environments capable of isolating high-level reasoning abilities of semantic maps from the complexities of noisy sensors and real-world observations. At the same time the survey establishes connections to SLAM-based approaches to underscore their methodological overlap and to bridge foundational techniques in robotics with emerging paradigms in embodied AI.
This survey is organized as follows. We begin the survey in Sec. 2 with background context to ground our discussion on semantic mapping. Sec. 3 introduces semantic maps and lays out the foundational components. We then explore different map structures in Sec. 4, followed by encoding strategies in Sec. 5. Sec. 6 reviews existing evaluation methodologies. In Sec. 7, we examine the key challenges in building semantic maps, leading into Sec. 8, which outlines promising directions for future research. We conclude in Sec. 9 by synthesizing the main insights and takeaways from the survey.
@@
Implicit (V) CMP (Gupta et al., 2017) MapNet (Henriques & Vedaldi, 2018) MultiON (Wani et al., 2020) RNR-Map (Kwon et al., 2023) SPTM (Savinov et al., 2018) NTS (Chaplot et al., 2020b) CMTP (Chen et al., 2021) VGM (Kwon et al., 2021) $\boxed{t}$ KPConv (Thomas et al., 2019) PointResNet (Desai et al., 2022) Neural Fusion (Mazur et al., 2023)
Implicit (VL) CoW (Gadre et al., 2023) VLMap (Huang et al., 2023a) Le-RNR-Map (Taioli et al., 2023) NLMap (Chen et al., 2023a) VLFM (Yokoyama et al., 2023) VoxPoser (Huang et al., 2023b) InstructNav (Long et al., 2024) OVL-MAP (Wen et al., 2025) LM-Nav (Shah et al., 2023) RoboHop (Garg et al., 2024) Clio (Maggio et al., 2024) OpenScene (Peng et al., 2023) CLIPFields (Shafiullah et al., 2023) ConceptFusion (Jatavallabhula et al., 2023) CLIP2Scene (Chen et al., 2023c) 3D aware ObjNav (Zhang et al., 2023a)
ConceptGraphs (Gu et al., 2023b) HOV-SG (Werby et al., 2024)
@@ 2 Background Reading
In this section, we provide background readings that help to contextualize the survey on semantic mapping. We first introduce key tasks (Sec. 2.1) in both robotics (Sec. 2.1.1) and embodied AI (Sec. 2.1.2), since semantic mapping is often studied in conjunction with such tasks. In Sec. 2.2, we discuss the basics (Sec. 2.2.1) and techniques of SLAM (Sec. 2.2.2) as well as the historic evolution of Semantic SLAM (Sec. 2.2.3), which has an overlap with the map building techniques that we later discuss throughout the survey. Next we briefly discuss two popular system design choices in Sec. 2.3, end-to-end learning (Sec. 2.3.1) and modular pipeline (Sec. 2.3.2), since it helps the readers understand how semantic map construction varies between end-to-end and modular approaches.
@@ 2.1.1 Robotics tasks
The era of modern robotics began with Unimate (Detesan & Moholea, 2024), the first industrial manipulator arm and Shakey (Center, 1984), the first mobile robot. Since then the field has evolved to have increased autonomy and complex reasoning skills. This has been enabled by research in various areas spanning basic navigation and obstacle avoidance to complex real-world capabilities involving perception, mapping, and manipulation. Early research focus on collision avoidance (Fox et al., 1997), Monte Carlo Localization (Thrun et al., 2001) and SLAM frameworks enabling robots to localize while mapping unknown environments (Thrun, 2002; Thrun & Montemerlo, 2006; Taheri & Xia, 2021). As sensors have improved over the years, semantic mapping gained traction, integrating object recognition and scene understanding into spatial maps (Salas-Moreno et al., 2013). Concurrently, robotic manipulation has become a central focus, with tasks such as pick-and-place, insertion, and rearrangement. Classical approaches like force-control-based manipulation (Raibert & Craig, 1981; Yoshikawa, 1985) have been complemented by modern reinforcement learning-based systems like QT-Opt (Kalashnikov et al., 2018) and Transporter Networks (Zeng et al., 2020). The field has also expanded to multi-task and long-horizon planning, where agents are required to complete sequences of interconnected tasks rather than isolated actions. A prominent framework in this area is Task and Motion Planning (TAMP) (Garrett et al., 2020), which integrates high-level symbolic reasoning (e.g., planning to "pick up the cup, open the door, and place the cup on the table") with low-level continuous motion planning (e.g., calculating joint angles and trajectories to execute those actions). This enables agents to operate in complex and dynamic environments. Recent trends include uncertainty-aware planning (Stachniss et al., 2005; Blanco et al., 2008a; Georgakis et al., 2022a; Carlone et al., 2013; Pan et al., 2019), semantic mapping (Rosinol et al., 2020b) and task planning in dynamic environments (Rosinol et al., 2020a). Another fast growing field in robotics is autonomous driving (Guan et al., 2024; Zhao et al., 2024), where recent trends increasingly rely on Bird’s Eye View (BEV) representations, which transform multi-view sensor inputs into a unified top-down map using end-to-end learned models. BEV is a common intermediate representation used to simplify reasoning about spatial layouts, obstacles, lanes and other semantic elements. Transformer-based architectures like BEVFormer (Li et al., 2024c) and BEVFusion (Liu et al., 2023c) have emerged as state-of-the-art, enabling spatially consistent, semantic-rich BEV maps that support tasks such as detection, planning and trajectory prediction. For example, LaneSegNet (Li et al., 2024b) trains an end-to-end system to predict lane segments from multi-view surrounding images. Unlike semantic maps that are discussed throughout this survey, BEV maps are typically short-lived local (per-frame) representations and specialized for driving scenarios. Overall, the advances in robotics collectively demonstrate a shift from reactive, single-task robots toward robust, general-purpose systems capable of operating autonomously in dynamic real-world environments.
@@ 2.2 Simultaneous Localization and Mapping (SLAM)
Although we briefly discuss robotics applications in Sec. 2.1, we address Simultaneous Localization and Mapping (SLAM) (Thrun et al., 1998a; 2000; Ferris et al., 2007; Huang et al., 2011; Carlone et al., 2024) separately here due to its close connection with semantic mapping, particularly the Semantic SLAM literature. While both aim to enrich spatial representations with semantic information, they differ in scope and purpose. Rooted in robotics, Semantic SLAM focuses on building globally consistent, pose-aware maps augmented with semantics to improve localization, mapping accuracy and long-term robustness. In contrast, semantic mapping in embodied AI emphasizes representations that support high-level reasoning and decision-making, often abstracting away precise localization and low-level sensor noise in favor of adaptability.
@@ 2.2.1 Basics of SLAM
where $\Omega_{t}$ is the information matrix (inverse covariance) that weights the measurements according to their uncertainty, i.e. more weight is given to a measurement that is very certain. $e_{t}^{T}(x)$ is the transpose of the error $e_{t}(x)$ and the term $e_{t}^{T}(x)~{}\Omega_{t}e_{t}(x)$ represents weighted squared error. In graph-based SLAM, this optimization corresponds to optimizing a factor graph (Loeliger, 2004; Dellaert et al., 2017), where nodes represent robot poses and edges represent spatial constraints, such that Equation 4 becomes:
$x^{*}=arg~{}\min_{x}\sum_{(i,j)\in\varepsilon}||z_{ij}-\hat{z}_{ij}(x_{i},x_{j})||^{2}_{\Omega_{ij}}$ (5)
where, $(i,j)\in\varepsilon$ represent edges in the pose graph, $z_{ij}$ is the relative pose measurement between nodes $i$ and $j$ , $\hat{z}_{ij}$ is the predicted relative measurement from current estimates, and ||.$||~{}.~{}||^{2}_{\Omega_{ij}}$ is the Mahalanobis distance (weighted squared error). This global optimization ensures that the estimated map and trajectory are as consistent as possible with the entire history of noisy sensor data.
@@ 2.2.2 SLAM Techniques
Active SLAM (Ahmed et al., 2023) enables an autonomous agent to actively choose its actions to improve its map and localization, rather than passively mapping the environment as it moves (Fig. 3-left). The SLAM system (Cadena et al., 2016; Pu et al., 2023) primarily consists of a front-end module and a back-end module (Fig. 3-right). The front-end module computes feature extraction, data association and feature classification on the sensor observations. Moreover it applies algorithms such as Iterative Closest Point (ICP) (He et al., 2017b) that aligns sensor observations from subsequent frames into consistent 3D geometry as well as loop closure (Tsintotas et al., 2022) that recognizes whether the current observation matches with a previously visited area. The backend module, on the other hand, is responsible for pose optimization and map estimation based on the data from the front-end. The back-end also provides feedback to the front-end for loop closure and verification. While some SLAM methods use LiDARs (Khan et al., 2021), others use camera as the primary sensor (visual SLAM or vSLAM), due to the availability of less expensive cameras and the advances in computer vision. Visual SLAM methods often extract and match geometric features (points, lines, or planes) (Mur-Artal et al., 2015; Yang & Scherer, 2017; Kaess, 2015; Cai et al., 2021) from the image, with the help of feature detection algorithms such as SIFT (Lowe, 2004), SURF (Bay et al., 2008), and ORB (Rublee et al., 2011). Others operate directly on the image pixel intensities (Newcombe et al., 2011b; Engel et al., 2014) and thus retain all information about the image. While these methods use RGB camera as the sensor, RGB-D SLAM (Newcombe et al., 2011a; Kaess et al., 2012; Endres et al., 2013; Whelan et al., 2016) uses RGB-D cameras to simultaneously collect color images as well as depth images. In contrast, Visual-Inertial SLAM (Mur-Artal & Tardós, 2017; Cheng et al., 2021) uses an additional IMU sensor to mitigate the effects of image blur and poor illumination from using camera sensor alone.
@@ 2.3 System design strategies
Designing systems for embodied agents involves a fundamental architectural choice between end-to-end learning and modular pipelines. While end-to-end approaches map raw sensory input directly to actions using a single neural network, modular systems decompose the task into interpretable components. Understanding this distinction is essential for situating semantic mapping within the broader system design, as it influences how maps are constructed, represented and utilized. This section briefly outlines these paradigms (Sec. 2.3.1 and Sec. 2.3.2) and discuss trade-offs between them in Sec. 2.3.3, to provide a background for the mapping methods discussed in the remainder of this survey.
@@ 2.3.1 End-to-end approaches
The embodied AI and robotics community has seen a lot of progress in training task-specific end-to-end models with reinforcement learning (RL) (Tai et al., 2017) that directly learns to predict discrete (Wani et al., 2020) or continuous actions (Kalapos et al., 2020) from visual observations (see Fig. 4). These methods may consist of an unstructured memory such as LSTM (Mei et al., 2016) (Dobrevski & Skočaj, 2021). However, such representation lacks reasoning about 3D space and geometry and fail to perform well in long-horizon path planning. This has lead to the development of approaches that build an intermediate map representation. Such a map can be implemented using differentiable operations so as to facilitate end-to-end training. Gupta et al. (2017) shows that an egocentric map built this way is beneficial for both PointNav and ObjectNav tasks, whereas Henriques & Vedaldi (2018) learns a global map for the task of localization. While these methods are trained using supervised learning, Wani et al. (2020) use RL to learn to predict actions based on an intermediate global map of the environment to address the complex MultiON task. In autonomous driving and planning, end-to-end approaches aim to learn a direct mapping from raw sensor inputs (e.g., images, LiDAR) to control outputs or motion trajectories. These methods often leverage transformer-based architectures (Lin et al., 2022) to jointly model perception, spatial reasoning and decision-making within a single framework. By optimizing the entire pipeline holistically, they promise improved robustness and efficiency, particularly in complex, dynamic environments. Recent work demonstrates that such end-to-end models can learn to implicitly reason about obstacles, goals, and traffic patterns. However challenges remain in interpretability, generalization and safety-critical validation. However, irrespective of the map representation or the mode of training, these approaches need to be retrained every time the task definition changes and even the basic skills required to perform a task need to be learned from scratch.

### v3

@@ 1 Introduction
To provide a principled understanding of semantic map-building methods, this survey categorizes the literature along two fundamental axes (see Fig.˜1): map structure (e.g., topological graphs, spatial grids, dense geometric, and hybrid representations) and semantic encoding (explicit annotations vs. learned implicit features). This taxonomy reflects core design choices that influence map scalability, interpretability, generalization, multi-modal fusion and querying. By organizing approaches in this manner, we aim to unify diverse strands of research (see Tab.˜1 for a summary of the reviewed papers), highlight the trade-offs between different representations and present key challenges and future opportunities in semantic mapping. This survey centers on semantic mapping approaches in the context of indoor mobile robots, a domain that offers a well-defined, practically relevant and technically rich environment for research. We focus on semantic mapping within embodied AI, that enables controlled environments capable of isolating high-level reasoning abilities of semantic maps from the complexities of noisy sensors and real-world observations. At the same time the survey establishes connections to SLAM-based approaches to underscore their methodological overlap and to bridge foundational techniques in robotics with emerging paradigms in embodied AI.
This survey is organized as follows. We begin the survey in Sec.˜2 with background context to ground our discussion on semantic mapping. Sec.˜3 introduces semantic maps and lays out the foundational components. We then explore different map structures in Sec.˜4, followed by encoding strategies in Sec.˜5. Sec.˜6 reviews existing evaluation methodologies. In Sec.˜7, we examine the key challenges in building semantic maps, leading into Sec.˜8, which outlines promising directions for future research. We conclude in Sec.˜9 by synthesizing the main insights and takeaways from the survey.
@@
Explicit (no semantics)
[…]
Topo SLAM (Choset & Nagatani, 2001)
Topomap (Blochliger et al., 2018)
Fast (Chen et al., 2022)
6D-SLAM (Nüchter et al., 2007)
Robust 3D mapping (May et al., 2009)
Droid-SLAM (Teed & Deng, 2021)
Voldor+ SLAM (Min & Dunn, 2021)
[…]
Explicit (semantics)
[…]
CM2 (Georgakis et al., 2022c)
[…]
Imitation (Duvallet et al., 2013)
Compact (Patki et al., 2019)
Multimodal (Arkin et al., 2020a)
[…]
ASHiTA (Chang et al., 2025)
@@
Following directions (Matuszek et al., 2010)
Conceptual (Zender et al., 2008)
SemanticGraph (Hemachandra et al., 2014)
Learning semantic maps (Walter et al., 2014)
Learning models (Hemachandra et al., 2015)
Inferring maps (Duvallet et al., 2016)
@@
LaneSegNet (Li et al., 2024b)
[…]
3DSG (Armeni et al., 2019)
[…]
Implicit to Explicit
[…]
SceneGraphFusion (Wu et al., 2021)
Implicit (V)
CMP (Gupta et al., 2017)
MapNet (Henriques & Vedaldi, 2018)
MultiON (Wani et al., 2020)
RNR-Map (Kwon et al., 2023)
SPTM (Savinov et al., 2018)
NTS (Chaplot et al., 2020b)
CMTP (Chen et al., 2021)
VGM (Kwon et al., 2021) $\boxed{t}$
KPConv (Thomas et al., 2019)
PointResNet (Desai et al., 2022)
Neural Fusion (Mazur et al., 2023)
Implicit (VL)
CoW (Gadre et al., 2023)
VLMap (Huang et al., 2023a)
Le-RNR-Map (Taioli et al., 2023)
NLMap (Chen et al., 2023a)
VLFM (Yokoyama et al., 2023)
VoxPoser (Huang et al., 2023b)
InstructNav (Long et al., 2024)
OVL-MAP (Wen et al., 2025)
LM-Nav (Shah et al., 2023)
RoboHop (Garg et al., 2024)
Clio (Maggio et al., 2024)
OpenScene (Peng et al., 2023)
CLIPFields (Shafiullah et al., 2023)
ConceptFusion (Jatavallabhula et al., 2023)
CLIP2Scene (Chen et al., 2023c)
3D aware ObjNav (Zhang et al., 2023a)
ConceptGraphs (Gu et al., 2024)
HOV-SG (Werby et al., 2024)
@@ 2 Background Reading
In this section, we provide background readings that help to contextualize the survey on semantic mapping. We first introduce key tasks (Sec.˜2.1) in both robotics (Sec.˜2.1.1) and embodied AI (Sec.˜2.1.2), since semantic mapping is often studied in conjunction with such tasks. In Sec.˜2.2, we discuss the basics (Sec.˜2.2.1) and techniques of SLAM (Sec.˜2.2.2) as well as the historic evolution of Semantic SLAM (Sec.˜2.2.3), which has an overlap with the map building techniques that we later discuss throughout the survey. Next we briefly discuss two popular system design choices in Sec.˜2.3, end-to-end learning (Sec.˜2.3.1) and modular pipeline (Sec.˜2.3.2), since it helps the readers understand how semantic map construction varies between end-to-end and modular approaches.
@@ 2.1.1 Robotics tasks
The era of modern robotics began with Unimate (Detesan & Moholea, 2024), the first industrial manipulator arm and Shakey (Center, 1984), the first mobile robot. Since then the field has evolved to have increased autonomy and complex reasoning skills. This has been enabled by research in various areas spanning basic navigation and obstacle avoidance to complex real-world capabilities involving perception, mapping, and manipulation. Early research focus on collision avoidance (Fox et al., 1997), Monte Carlo Localization (Thrun et al., 2001) and SLAM frameworks enabling robots to localize while mapping unknown environments (Thrun, 2002; Thrun & Montemerlo, 2006; Taheri & Xia, 2021). As sensors have improved over the years, semantic mapping gained traction, integrating object recognition and scene understanding into spatial maps (Salas-Moreno et al., 2013). Concurrently, robotic manipulation has become a central focus, with tasks such as pick-and-place, insertion, and rearrangement. Classical approaches like force-control-based manipulation (Raibert & Craig, 1981; Yoshikawa, 1985) have been complemented by modern reinforcement learning-based systems like QT-Opt (Kalashnikov et al., 2018) and Transporter Networks (Zeng et al., 2020). The field has also expanded to multi-task and long-horizon planning, where agents are required to complete sequences of interconnected tasks rather than isolated actions. A prominent framework in this area is Task and Motion Planning (TAMP) (Garrett et al., 2020), which integrates high-level symbolic reasoning (e.g., planning to "pick up the cup, open the door, and place the cup on the table") with low-level continuous motion planning (e.g., calculating joint angles and trajectories to execute those actions). This enables agents to operate in complex and dynamic environments. Recent trends include uncertainty-aware planning (Stachniss et al., 2005; Blanco et al., 2008a; Georgakis et al., 2022a; Carlone et al., 2013; Pan et al., 2019), semantic mapping (Rosinol et al., 2020b) and task planning in dynamic environments (Rosinol et al., 2020a). Another fast growing field in robotics is autonomous driving (Guan et al., 2024; Zhao et al., 2024), where recent trends increasingly rely on Bird’s Eye View (BEV) representations, which transform multi-view sensor inputs into a unified top-down map using end-to-end learned models. BEV is a common intermediate representation used to simplify reasoning about spatial layouts, obstacles, lanes and other semantic elements. Transformer-based architectures like BEVFormer (Li et al., 2024c) and BEVFusion (Liu et al., 2023b) have emerged as state-of-the-art, enabling spatially consistent, semantic-rich BEV maps that support tasks such as detection, planning and trajectory prediction. For example, LaneSegNet (Li et al., 2024b) trains an end-to-end system to predict lane segments from multi-view surrounding images. Unlike semantic maps that are discussed throughout this survey, BEV maps are typically short-lived local (per-frame) representations and specialized for driving scenarios. Overall, the advances in robotics collectively demonstrate a shift from reactive, single-task robots toward robust, general-purpose systems capable of operating autonomously in dynamic real-world environments.
@@ 2.2 Simultaneous Localization and Mapping (SLAM)
Although we briefly discuss robotics applications in Sec.˜2.1, we address Simultaneous Localization and Mapping (SLAM) (Thrun et al., 1998a; 2000; Ferris et al., 2007; Huang et al., 2011; Carlone et al., 2024) separately here due to its close connection with semantic mapping, particularly the Semantic SLAM literature. While both aim to enrich spatial representations with semantic information, they differ in scope and purpose. Rooted in robotics, Semantic SLAM focuses on building globally consistent, pose-aware maps augmented with semantics to improve localization, mapping accuracy and long-term robustness. In contrast, semantic mapping in embodied AI emphasizes representations that support high-level reasoning and decision-making, often abstracting away precise localization and low-level sensor noise in favor of adaptability.
@@ 2.2.1 Basics of SLAM
where $\Omega_{t}$ is the information matrix (inverse covariance) that weights the measurements according to their uncertainty, i.e. more weight is given to a measurement that is very certain. $e_{t}^{T}(x)$ is the transpose of the error $e_{t}(x)$ and the term $e_{t}^{T}(x)~\Omega_{t}e_{t}(x)$ represents weighted squared error. In graph-based SLAM, this optimization corresponds to optimizing a factor graph (Loeliger, 2004; Dellaert et al., 2017), where nodes represent robot poses and edges represent spatial constraints, such that Equation˜4 becomes:
[…]
where, $(i,j)\in\varepsilon$ represent edges in the pose graph, $z_{ij}$ is the relative pose measurement between nodes $i$ and $j$ , $\hat{z}_{ij}$ is the predicted relative measurement from current estimates, and $||~.~||^{2}_{\Omega_{ij}}$ is the Mahalanobis distance (weighted squared error). This global optimization ensures that the estimated map and trajectory are as consistent as possible with the entire history of noisy sensor data.
@@ 2.2.2 SLAM Techniques
Active SLAM (Ahmed et al., 2023) enables an autonomous agent to actively choose its actions to improve its map and localization, rather than passively mapping the environment as it moves (Fig.˜3-left). The SLAM system (Cadena et al., 2016; Pu et al., 2023) primarily consists of a front-end module and a back-end module (Fig.˜3-right). The front-end module computes feature extraction, data association and feature classification on the sensor observations. Moreover it applies algorithms such as Iterative Closest Point (ICP) (He et al., 2017b) that aligns sensor observations from subsequent frames into consistent 3D geometry as well as loop closure (Tsintotas et al., 2022) that recognizes whether the current observation matches with a previously visited area. The backend module, on the other hand, is responsible for pose optimization and map estimation based on the data from the front-end. The back-end also provides feedback to the front-end for loop closure and verification. While some SLAM methods use LiDARs (Khan et al., 2021), others use camera as the primary sensor (visual SLAM or vSLAM), due to the availability of less expensive cameras and the advances in computer vision. Visual SLAM methods often extract and match geometric features (points, lines, or planes) (Mur-Artal et al., 2015; Yang & Scherer, 2017; Kaess, 2015; Cai et al., 2021) from the image, with the help of feature detection algorithms such as SIFT (Lowe, 2004), SURF (Bay et al., 2008), and ORB (Rublee et al., 2011). Others operate directly on the image pixel intensities (Newcombe et al., 2011b; Engel et al., 2014) and thus retain all information about the image. While these methods use RGB camera as the sensor, RGB-D SLAM (Newcombe et al., 2011a; Kaess et al., 2012; Endres et al., 2013; Whelan et al., 2016) uses RGB-D cameras to simultaneously collect color images as well as depth images. In contrast, Visual-Inertial SLAM (Mur-Artal & Tardós, 2017; Cheng et al., 2021) uses an additional IMU sensor to mitigate the effects of image blur and poor illumination from using camera sensor alone.
@@ 2.3 System design strategies
Designing systems for embodied agents involves a fundamental architectural choice between end-to-end learning and modular pipelines. While end-to-end approaches map raw sensory input directly to actions using a single neural network, modular systems decompose the task into interpretable components. Understanding this distinction is essential for situating semantic mapping within the broader system design, as it influences how maps are constructed, represented and utilized. This section briefly outlines these paradigms (Sec.˜2.3.1 and Sec.˜2.3.2) and discuss trade-offs between them in Sec.˜2.3.3, to provide a background for the mapping methods discussed in the remainder of this survey.
@@ 2.3.1 End-to-end approaches
The embodied AI and robotics community has seen a lot of progress in training task-specific end-to-end models with reinforcement learning (RL) (Tai et al., 2017) that directly learns to predict discrete (Wani et al., 2020) or continuous actions (Kalapos et al., 2020) from visual observations (see Fig.˜4). These methods may consist of an unstructured memory such as LSTM (Mei et al., 2016) (Dobrevski & Skočaj, 2021). However, such representation lacks reasoning about 3D space and geometry and fail to perform well in long-horizon path planning. This has lead to the development of approaches that build an intermediate map representation. Such a map can be implemented using differentiable operations so as to facilitate end-to-end training. Gupta et al. (2017) shows that an egocentric map built this way is beneficial for both PointNav and ObjectNav tasks, whereas Henriques & Vedaldi (2018) learns a global map for the task of localization. While these methods are trained using supervised learning, Wani et al. (2020) use RL to learn to predict actions based on an intermediate global map of the environment to address the complex MultiON task. In autonomous driving and planning, end-to-end approaches aim to learn a direct mapping from raw sensor inputs (e.g., images, LiDAR) to control outputs or motion trajectories. These methods often leverage transformer-based architectures (Lin et al., 2022) to jointly model perception, spatial reasoning and decision-making within a single framework. By optimizing the entire pipeline holistically, they promise improved robustness and efficiency, particularly in complex, dynamic environments. Recent work demonstrates that such end-to-end models can learn to implicitly reason about obstacles, goals, and traffic patterns. However challenges remain in interpretability, generalization and safety-critical validation. However, irrespective of the map representation or the mode of training, these approaches need to be retrained every time the task definition changes and even the basic skills required to perform a task need to be learned from scratch.